# Model Provenance Testing for Large Language Models

### v1

@@ 6 Conclusion
Our work formulates the model provenance testing problem for large language models which has many applications such as in detection of misuse of licensing and terms of use or problematic customized models. We present an approach based on statistical testing with minimal assumptions that has high accuracy for real-world benchmarks. Our key insight is that models derived through standard customization approaches maintain a level of similarity to their parent model that is statistically distinguishable from unrelated parents. We evaluate this observation empirically, together with our approach and several optimizations in Section 4. We find our proposed method to be practical for deciding LLM provenance.
@@ Acknowledgments
We thank Xu Louise, Bo Wang, Jason Zhijingcheng Yu, and Mallika Prabhakar for their valuable feedback on the draft. This research was supported by the Crystal Centre at National University of Singapore.
Appendix A Models and Benchmarks
@@
The first benchmark, called Bench-A, consists of LLM pairs for model provenance constructed from popular pre-trained models and their fine-tuned derivatives. To build this benchmark, we manually selected 10 widely-used pre-trained models (refer to Tbl. 3) with between 1 billion and 4 billion parameters (the upper bound was determined by our GPU memory constraints). Among these, we purposefully included four pairs of architecturally similar models from Meta, Microsoft, Google, and AliBaba to evaluate our tester’s ability to distinguish between closely related base models and to have some control models. For each pre-trained model, we then randomly sampled 10 fine-tuned derivatives using the Hugging Face API (i.e. use highly reliable ground truth verification), prioritizing diversity in model creators. This sampling strategy resulted in 100 derived models, that constitute Bench-A.
Table 3: All 10 pre-trained LLMs from Bench-A.
@@
Appendix B Sampling Prompts
@@ 5 The phonetics lecturer explained the intricacies of the
Appendix C Additional Tables
@@
Table 7: Most similar pre-trained models from Bench-B sorted for $k=1$ (no advanced prompt sampling), and their corresponding values for $k=4,16,64$.
@@
Appendix D Testing with Known Parent
We evaluate the provenance tester for cases where the parent model is given, addressing whether a pair of models $(P,C)$ constitutes a provenance pair. We construct test pairs from both Bench-A and Bench-B benchmarks. From Bench-A, we take all 100 true pairs $(P_{i},C_{i})$ and create 100 false pairs $(\tilde{P_{i}},C_{i})$ by selecting one random non-parent $\tilde{P_{i}}\neq P_{i}$ for each child $C_{i}$. This ensures a balanced dataset where random guessing would achieve 50% accuracy. We similarly obtain 766 testing pairs from Bench-B  [^7] Unlike Bench-A, Bench-B already contains (child) models that have no known parent among the 57 pre-trained models, and we use these as one of the negative pairs..
Results from testing both benchmarks are shown in Figures 6 and 7. As expected, the tester performs better when the suspected parent is known compared to cases with unspecified parents. While recall remains unchanged, precision reaches 100%. This improvement occurs because false positives now require both that the statistical hypothesis test returns the wrong parent and that this wrong parent matches the suspected parent. The precision with known parents ($P_{knownparent}$) is lower bounded by the precision with unspecified parents ($P_{unspecified}$), and depends on how we sample the incorrect parents. With uniform random sampling, it can be estimated as $P_{knownparent}=100-\frac{100-P_{unspecified}}{n-1}$, where $n$ is the number of pre-trained models. For both benchmarks, this yields precision slightly below 100%.
Figure 6: Precision and recall of running the model provenance tester on Bench-A without known parent (dark blue and dark red), and with a suspected parent (light blue and light red).
Figure 7: Precision and recall of running the model provenance tester on Bench-B without known parent (dark blue and dark red), and with a suspected parent (light blue and light red).
Appendix E Advanced Sampling of Prompts
@@
$Prompts\leftarrow\emptyset$triangleright Initialize empty set of prompts
$H\leftarrow F\cup C$triangleright All models
$\text{same}[i][j]\leftarrow 0\text{ for all }i,j\in[|H|]$triangleright Counter of times two models produced the same output token
@@
     $x_{1},\ldots,x_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample $k$ prompts
     $Score[j]\leftarrow 0\text{ for all }j\leftarrow 1$ to $k$triangleright Scores across $k$ prompts
@@
$l\leftarrow\operatorname*{arg\,max}_{j}\text{Score}[j]$triangleright Find largest score
$Prompts\leftarrow Prompts\cup\{x_{l}\}$triangleright Add that prompt
return $Prompts$
Appendix F Reducing Offline Queries with Best Arm Identification
@@
Model $g$, candidate set $F=\{f_{1},\ldots,f_{s}\}$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T$, significance parameter $\alpha$, maximum average prompts per model $N$
$M\leftarrow F\cup C$triangleright Set of all models
$U(t,\alpha):=\sqrt{\frac{\log(4t^{2}/\alpha)}{2t}}$triangleright Confidence interval for BAI
$\text{hits}[m]\leftarrow 0\text{ for all }m\in M$triangleright # same tokens with $g$
$\text{tots}[m]\leftarrow 0\text{ for all }m\in M$triangleright # queried
$A\leftarrow M$triangleright Active set of models
@@
     $x\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$triangleright Sample prompt
     $y_{g}\leftarrow g(x)$triangleright Query $g$
@@
         $y_{m}\leftarrow m(x)$triangleright query model $m$
         $hits[m]{+}=\mathds{1}(y_{g},y_{m})$triangleright Update hit counters
         $tots[m]{+}=1$triangleright Update total queries
@@
     $\mu_{best}\leftarrow\max_{m\in M}\{\frac{\text{hits}[m]}{\text{tots}[m]}\}$triangleright Find best $\mu$
     $u\leftarrow U(t,\alpha)$triangleright Confidence radius
@@
              $A\leftarrow A\setminus\{m\}$triangleright $\mu_{m}$ too far from best

### v2

@@ 5 Conclusion
Our work formulates the model provenance testing problem for LLMs which has many applications such as in detection of misuse of terms of use or vulnerable customized models. We present an approach based on statistical testing with minimal assumptions that has high accuracy for real-world benchmarks. Our key insight is that models derived through standard customization approaches maintain a level of similarity to their parent model that is statistically distinguishable from unrelated parents. We evaluate this observation empirically, together with our approach and several optimizations. We find our proposed method to be practical for deciding LLM provenance.
@@ Acknowledgments
We thank Xu Louise, Bo Wang, Jason Zhijingcheng Yu, and Mallika Prabhakar for their valuable feedback on the draft. This research is supported by a Singapore Ministry of Education (MOE) Tier 2 grant MOE-T2EP20124-0007. This material is based upon work supported by the U.S. National Science Foundation under Grant No. 2531010. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF.
Appendix A Related Work
We further detail prior work’s required access to the model and any additional knowledge to identify or show ownership of a model, in particular large language models in Table 3.
White-box setting is the most prevalent in existing works. In particular, it is assumed that the analysis is done by the model owner or defender who has access to its model weights. Our approach is fully black-box based and it only analyzes the first output token of each query. For instance, in [53], the defender has access to their own model’s parameters and the last linear layer as a fingerprint, and aims to verify ownership through API access to the suspect model. LLMmap considers only different versions of the same model and 5 model variants that have resulted from fine-tuning [34]. LLMmap proposes to train a model on a set of queries to detect unusual responses from a model which may come in response to an attacker’s fingerprinting probing. It uses a pre-trained textual embedding model to generate vector representations of the query and responses that they then use to classify and identify models. It also uses heuristics to identify the most promising prompts for model identification. Other LLM fingerprinting approaches uses system-level statistics to analyze the fingerprint models such as inter-token times when these are available either locally or over the network [1].
Table 3: Comparison of our work to related work based on fingerprinting.
Work Access Type Technique Additional Knowledge
Our Black-box with only the next token Hypothesis testing with statistical guarantees No
[34] Black-box with logits Proposes to train a model on a set of queries Pre-trained embeddings for classifier; Queries manual crafted and domain knowledge
[57] White-box model internal representations Classifiers trained on representations No
[51] White-box. Models are embedded with fingerprint key pairs Embedding robust fingerprints into the owner model. Changes in training procedure
[1] System-level statistics such as inter-token times Deep learning to classify Network traffic
[4] Same as LLMmap LLMmap and other dynamic fingerprinting Judge model, additional access to classifier models and manual fingerprinting
[53] White-box. Model’s parameters and the last linear layer as a fingerprint and API access to the suspect model. Vector similarity No
[29] Takes datasets of AI-generated texts and classifies them in families Linguistic features for similar styles No
[54] White-box access to base model, black-box access to the suspect model A proactive fine-tuning approach to modify the output of the model No
[28] Knowledge of the training dataset Statistical testing and distance to margin Classification models
[30] Log-probabilities of the target and random perturbations of the input passage to another model Used to decide machine-generated text Additional model
Appendix B Basic Tester
Algorithm 2 Model Provenance Tester for Pair $(f,g)$
$x_{1},\ldots,x_{T}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$ $\triangleright$ Sample T prompts
$\mu\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(f(x_{j})=g(x_{j}))$ $\triangleright$ Calc sim of f and g
for $i\leftarrow 1$ to $m$ do
  $\mu_{i}\leftarrow\frac{1}{T}\sum_{j=1}^{T}\mathds{1}(c_{i}(x_{j})=g(x_{j}))$ $\triangleright$ Calc sim of $c_{i}$ and $g$
  $p_{i}\leftarrow\text{ZTest}(\mu,\mu_{i},T)$ $\triangleright$ Obtain p-values
end for
$(p_{(1)},\ldots,p_{(m)})\leftarrow\text{Sort}(p_{1},\ldots,p_{m})$ $\triangleright$ Sort p-values
for $k\leftarrow 1$ to $m$ do
  $\alpha_{k}\leftarrow\alpha/(m-k+1)$ $\triangleright$ Holm-Bonferroni adjustment
  end if
end forreturn True $\triangleright$ Is a provenance pair
A pseudo-code of the basic tester is given in Algorithm 2. In summary, when the procedure returns True, it ensures that the total family-wise error rate (FWER) is controlled at the significance level of $\alpha=0.05$ or lower. This means we can confidently state that the similarity between $f$ and $g$ is significantly higher than between $g$ and any control model, supporting the existence of a provenance relationship. Conversely, when the procedure returns False, it indicates that we could not establish this higher similarity with the desired level of statistical significance. This may occur either because there is genuinely no significant similarity indicative of provenance or because the test lacked sufficient power under the given parameters (e.g., sample size or number of prompts) to detect it.
Appendix C Models and Benchmarks
@@
The first benchmark, called Bench-A, consists of LLM pairs for model provenance constructed from popular pre-trained models and their fine-tuned derivatives. To build this benchmark, we manually selected 10 widely-used pre-trained models (refer to Tbl. 4) with between 1 billion and 4 billion parameters (the upper bound was determined by our GPU memory constraints). Among these, we purposefully included four pairs of architecturally similar models from Meta, Microsoft, Google, and AliBaba to evaluate our tester’s ability to distinguish between closely related base models and to have some control models. For each pre-trained model, we then randomly sampled 10 fine-tuned derivatives using the Hugging Face API (i.e. use highly reliable ground truth verification), prioritizing diversity in model creators. This sampling strategy resulted in 100 derived models, that constitute Bench-A.
Table 4: All 10 pre-trained LLMs from Bench-A (left) and top 10 from Bench-B (right).
@@
Examples of Candidates’ Domains from Bench-A/B.  We find examples of candidates for domains such as financial, biology and medical among others. See Table 5 for a list of examples.
Table 5: Domain-specific Hugging Face models included in our analysis.
Domain Model ID Hugging Face URL
Medical medical_transcription_generator https://huggingface.co/alibidaran/medical_transcription_generator
Medical gpt2-large-medical https://huggingface.co/Locutusque/gpt2-large-medical
Medical healthbot https://huggingface.co/Anjanams04/healthbot
Law bloom-560m-finetuned-fraud https://huggingface.co/jslin09/bloom-560m-finetuned-fraud
Financial FinguAI-Chat-v1 https://huggingface.co/FINGU-AI/FinguAI-Chat-v1
Financial FinOPT-Lincoln https://huggingface.co/MayaPH/FinOPT-Lincoln
Biology distilgpt2-finetuned-microbiology https://huggingface.co/as-cle-bert/distilgpt2-finetuned-microbiology
Astrology astroGPT https://huggingface.co/stevhliu/astroGPT
Game magic-the-gathering https://huggingface.co/minimaxir/magic-the-gathering
Game chessgpt2-medium-l https://huggingface.co/dakwi/chessgpt2-medium-l
Experimental Setup.
Appendix D Sampling Prompts
@@ 5 The phonetics lecturer explained the intricacies of the
Appendix E The Effectiveness of Rejection sampling
Certain pre-trained models from Bench-A and Bench-B exhibit a high degree of similarity when comparing their output tokens generated from random prompts. Table 8 presents the top $5$ most similar model pairs from Bench-B, measured by the percentage of matching output tokens when tested on $1,000$ random prompts.
@@ Appendix E The Effectiveness of Rejection sampling
Table 8: Most similar pre-trained models from Bench-B sorted for $k=1$ and their corresponding values for $k=4,16,64$.
@@
Appendix F Testing with Known Parent
[Uncaptioned image][Uncaptioned image]
Figure 6: Precision/recall the tester on Bench-A (left) and Bench-B (right) without known parent (dark blue and dark red), and with a suspected parent (light blue and light red) ).
Figure 7: Precision and recall of running the model provenance tester on Bench-A with different significance levels $\alpha$.
We evaluate the provenance tester for cases where the parent model is given, addressing whether a pair of models $(P,C)$ constitutes a provenance pair.
We construct test pairs from both Bench-A and Bench-B benchmarks. From Bench-A, we take all 100 true pairs $(P_{i},C_{i})$ and create 100 false pairs $(\tilde{P_{i}},C_{i})$ by selecting one random non-parent $\tilde{P_{i}}\neq P_{i}$ for each child $C_{i}$. This ensures a balanced dataset where random guessing would achieve 50% accuracy. We similarly obtain 766 testing pairs from Bench-B  [^5] Unlike Bench-A, Bench-B already contains (child) models that have no known parent among the 57 pre-trained models, and we use these as one of the negative pairs..
Results from testing both benchmarks are shown in Figure 6. As expected, the tester performs better when the suspected parent is known compared to cases with unspecified parents. While recall remains unchanged, precision reaches 100%. This improvement occurs because false positives now require both that the statistical hypothesis test returns the wrong parent and that this wrong parent matches the suspected parent. The precision with known parents ($P_{knownparent}$) is lower bounded by the precision with unspecified parents ($P_{unspecified}$), and depends on how we sample the incorrect parents. With uniform random sampling, it can be estimated as $P_{knownparent}=100-\frac{100-P_{unspecified}}{n-1}$, where $n$ is the number of pre-trained models. For both benchmarks, this yields precision slightly below 100%.
Appendix G Precision/Recall for Different $\alpha$
We also run the model provenance tester for Bench-A with different significance level $\alpha\in\{0.01,0.05,0.1\}$. The results of these experiments are given in Figure 7. As evident from the results, it is clear that the significance level does not have any major impact on the efficiency of the tester.
Appendix H Advanced Sampling of Prompts
@@
$Prompts\leftarrow\emptyset$ $\triangleright$ Initialize empty set of prompts
$H\leftarrow F\cup C$ $\triangleright$ All models
$\text{same}[i][j]\leftarrow 0\text{ for all }i,j\in[|H|]$ $\triangleright$ Counter of times two models produced the same output token
@@
  $x_{1},\ldots,x_{k}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$ $\triangleright$ Sample $k$ prompts
  $Score[j]\leftarrow 0\text{ for all }j\leftarrow 1$ to $k$ $\triangleright$ Scores across $k$ prompts
@@
$l\leftarrow\operatorname*{arg\,max}_{j}\text{Score}[j]$ $\triangleright$ Find largest score
$Prompts\leftarrow Prompts\cup\{x_{l}\}$ $\triangleright$ Add that prompt
return PromptsPrompts
Appendix I Reducing Offline Queries with Best Arm Identification
@@
Model $g$, candidate set $F=\{f_{1},\ldots,f_{s}\}$, set of control models $C=\{c_{1},\ldots,c_{m}\}$, prompt space $\Omega$, number of prompts $T$, significance parameter $\alpha$, maximum average prompts per model NN
$M\leftarrow F\cup C$ $\triangleright$ Set of all models
$U(t,\alpha):=\sqrt{\frac{\log(4t^{2}/\alpha)}{2t}}$ $\triangleright$ Confidence interval for BAI
$\text{hits}[m]\leftarrow 0\text{ for all }m\in M$ $\triangleright$ # same tokens with $g$
$\text{tots}[m]\leftarrow 0\text{ for all }m\in M$ $\triangleright$ # queried
$A\leftarrow M$ $\triangleright$ Active set of models
@@
  $x\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\Omega$ $\triangleright$ Sample prompt
  $y_{g}\leftarrow g(x)$ $\triangleright$ Query $g$
@@
   $y_{m}\leftarrow m(x)$ $\triangleright$ query model $m$
   $hits[m]{+}=\mathds{1}(y_{g},y_{m})$ $\triangleright$ Update hit counters
   tots[m]+=1tots[m]{+}=1 $\triangleright$ Update total queries
@@
  $\mu_{best}\leftarrow\max_{m\in M}\{\frac{\text{hits}[m]}{\text{tots}[m]}\}$ $\triangleright$ Find best $\mu$
  $u\leftarrow U(t,\alpha)$ $\triangleright$ Confidence radius
@@
     $A\leftarrow A\setminus\{m\}$ $\triangleright$ $\mu_{m}$ too far from best